OpenStudy (anonymous):

A particle is moving with the given data. Find the position of the particle. a(t)=10sin(t)+3cos(t) s(0)=0 s(2pi)=12

zepdrix (zepdrix):

So you're getting into anti-differentiation now i see? :) They gave us a function for the particles ACCELERATION. Acceleration is the derivative of Velocity. Velocity is the derivative of Position. So we'll need to anti-differentiate 2 times. If we anti-differentiate a(t) we'll get,\[\huge v(t)=-10\cos(t)+3\sin(t)+c\]Understand how the trig functions changed? Also, whenever we anti-differentiate an unknown constant value will be tacked onto the end of the problem. If you need an explanation as to why the +c is there, let me know. Ok where you stuck at? :)

OpenStudy (anonymous):

Okay, I understand that (:

OpenStudy (anonymous):

Now, do I need to find the anti derivative of that anti derivative..?

zepdrix (zepdrix):

Yes c:

OpenStudy (anonymous):

hmm, can I hire you to teach me? (; you're so much easier to understand then my chinese professor hahaha!

zepdrix (zepdrix):

chinese? oh boy that sounds rough :3 lol

OpenStudy (anonymous):

hahah oh it is rough. She's insane in the membrane.

zepdrix (zepdrix):

heh :D do you understand how to find the next anti-derivative? it's a little trickier than the first, since something will happen with our c term.

OpenStudy (anonymous):

oh no! Let me try (:

OpenStudy (anonymous):

I got... -10sin(t)-3cos(t).... now im stuck?

zepdrix (zepdrix):

When we anti-differentiate a second time, this is what we should end up with, \[\huge s(t)=-10\sin(t)-3\cos(t)+cx+d\]

OpenStudy (anonymous):

uhm, why cx+d?

zepdrix (zepdrix):

So think about this a second. When you take a DERIVATIVE, every term is DECREASING in power by 1 degree yes? (at least for the powers of x). So if we had something like \[\huge \frac{d}{dx}3x^1 \qquad \rightarrow \qquad 3x^0 \qquad \rightarrow \qquad 3\] Now let's think about that in reverse for a moment... If I start with a constant, I can think of it as a constant times x^0. (Because x^0 is just 1). From there, we want to INCREASE the power of X, since we're anti-differentiating.

OpenStudy (anonymous):

ohh... I'm getting ya...

zepdrix (zepdrix):

\[3 \qquad \rightarrow \qquad 3x^0\]Anti-differentiating gives us,\[3x^1\]

zepdrix (zepdrix):

In our case, it's not a 3, it's a C, but the principle still applies. As for the d, that is just a NEW constant term that appeared! I could have called it C2 or something perhaps, but D seemed fine... Remember we get a constant EACH TIME we anti-differentiate.

OpenStudy (anonymous):

Okay so now.. what are s(0)=0 and s(2pi)=12 for?

zepdrix (zepdrix):

Those are specific values of s(t) that they provided us with. What we have come up with so far, is a GENERAL SOLUTION for s(t). It includes a couple of UNKNOWN constants that we can't solve for unless we're given initial conditions. Well, we were! :) So we'll use those values to solve for c and d.

zepdrix (zepdrix):

\[\large s(0)=0 \qquad \rightarrow \qquad 0=-10\sin(0)-3\cos(0)+3(0)+d\]

zepdrix (zepdrix):

Understand how we did that? We set it equal to 0, before s(0) is equal to zero. And then plug in 0 for all of our x values. We should be able to easily solve for d at this point.

OpenStudy (anonymous):

d=3?

zepdrix (zepdrix):

k sounds good c:

zepdrix (zepdrix):

Then use the other given value of s(t) to find c \:D/

OpenStudy (anonymous):

s(2pi)=12... 12=-10sin(2pi)-3cos(2pi)+3c+3?

OpenStudy (anonymous):

did i set that up wrong?

zepdrix (zepdrix):

all looks good except the c part. We have already solved for D, so D is 3. But +cx should change to C(2pi) right?

zepdrix (zepdrix):

i dunno why i changed from lowercase c to uppercase, that wasn't intentional lol

OpenStudy (anonymous):

ohhh okay (:

OpenStudy (anonymous):

I simplified it down to... 2pi*c=12

OpenStudy (anonymous):

6/pi=c yay! I got it (:

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

hha thanks so much~!